OpenStudy (anonymous):

Ben bowled 145 and 182 in his first two games. What must he bowl in his third game to have an average of at least 170?

OpenStudy (btaylor):

(145+182+X)/3=170 170 x 3 = 145+182+X 510 = 327+X X=183 Ben needs a 183.